don williams uses his small boat to go 3 miles up stream. against the current, the trip takes 3/5 hoours. with the current the trip takes 1/5 hours how fast is the boat on still water? what is the speed of the current?
Answer by mananth(12270)

boat speed x mph
current speed = y
...
upstream speed = x-y
downstream speed x+y
..
distance = 3 miles
..
t=d/r
..
3/(x-y)=3/5
cross multiply
3(x-y)15
/3
x-y =5.........................1
..
3/(x+y)=1/5
x+y =15.........................2
..
add 1 & 2
2x=20
/2
x= 10.
speed of boat in still water = 10 mph.
speed of current = 5 mph
m.ananth@hotmail.ca